For each of the following numbers, find the smallest whole number by which it should be multiplied so as to get a perfect square number. Also find the square root of the square number so obtained. 
 
768

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

768 = 2×2×2×2×2×2×2×2×3
= (2×2)×(2×2)×(2×2)×(2×2)×3
Here, 3 cannot be paired.
∴ We will multiply 768 by 3 to get perfect square.New number = 768×3 = 2304

2304 = 2×2×2×2×2×2×2×2×3×3
⇒2304 = (2×2)×(2×2)×(2×2)×(2×2)×(3×3)
2304 = 22×22×22×22×32
⇒ 2304 = (2×2×2×2×3)2
⇒ √2304 = 2×2×2×2×3 = 48
Thanks